2. A latte in the U.S. costs $4 and in a Central American country it costs 28 pesos. The exchange rate between these two currencies is 8 pesos per dollar. Based on this information the implied exchange rate is relatively cheaper/more expensive (circle one) in the Central American country. If GDP per capita is 64,000 pesos per dollar then this converts to S official exchange rate. The purchasing power parity adjusted value of GDP per capita would be $ pesos per dollar and goods are using the

Need help, please show work
Answer #1


Cost of latte in US = $4

Cost of latte in Central American Country = 28 pesos

Implied exchange rate = 28/4 = 7 pesos per dollar

Officially 8 pesos can be exchanged for $1 but in terms of goods, goods worth 7 pesos can be exchanged for $1.

This means price level is higher in Central American Country relative to the United States.

So,

Based on this information, the implied exchange rate is 7 pesos per dollar and goods are more expensive in the Central American Country.

Official exchange rate = 8 pesos per dollar

Converting pesos into dollars -

64,000 pesos = 64,000/8 = $8,000

So,

If GDP per capita is 64,000 pesos per dollar then this convert to $8,000 using the official exchange rate.

Implied exchange rate = 7 pesos per dollar

Converting pesos into dollars -

64,000 pesos = 64,000/7 = $9,142.85

So,

The purchasing power parity adjusted value of GDP per capita would be $9,142.85